OpenStudy (anonymous):

(Please help!) One way to describe the concentration of an acid is as a percent by volume. For example, in 40 mL of a 30% acid solution, the volume of pure acid is 40 x 30/100 or 12 mL, and the volume of water is 40-12 or 28 mL. If 5 mL of pure acid is mixed with 20 mL of water to give 25 mL of acid solution, the concentration of the solution is given by 5/25 x 100% = 20 %. If water is mixed with 50 mL of 40% acid solution, write an equation that describe the acid concentration, C(x), as a function of the volume of water added, x.

OpenStudy (anonymous):

The answer is C(x) = (20/50 + x)*100% and I have to graph C(x) versus x. But how do I graph it on the graphing calculator?

OpenStudy (anonymous):

The answer on the back of the book is: 

hero (hero):

If that's the answer, then just graph this : C(x) = (20/50 + x) forget about the 100% because 100% = 1

OpenStudy (anonymous):

but on my graphing calculator it's a diagonal line.. o.o

OpenStudy (anonymous):

it's still a diagonal squiggly line.. o.o

hero (hero):

Are you sure you have posted the proper function?

OpenStudy (anonymous):

ohh.. I plot that on my graph and yes I posted the proper function D:

hero (hero):

Then why are there negative numbers associated with it? Something has to be wrong

OpenStudy (anonymous):

Okay so the function of the volume of water added, x is: \[C(x) = 20/50 + x \times 100\]

OpenStudy (anonymous):

*100%

hero (hero):

Your function isn't clear enough because it could be interpreted in two different ways

OpenStudy (anonymous):

oh boy. D: So with one of the two ways, what function will have the same graph as i drew above?

hero (hero):

\[C(x) = \frac{20}{50} + x \times 100\] or \[C(x) = \frac{20}{50+x} \times 100\]

hero (hero):

And I forgot the percent sign.

OpenStudy (anonymous):

ohh.. it shows from the back of my book the second function..

hero (hero):

idk, there is something about the function you must be misinterpreting

OpenStudy (anonymous):

in the question above, what function would you get?

hero (hero):

Okay, well in that case, it is : \[C(x) = \frac{2000}{50+x}\]

hero (hero):

Graph that and tell me what you get

OpenStudy (anonymous):

I get a diagonal line.. o.o

hero (hero):

The something is wrong with your calc or you're just not graphing correctly

OpenStudy (anonymous):

how do I graph it correctly?

hero (hero):

Enter this exactly: 2000/(50+x)

OpenStudy (mertsj):

If you have 50 ml of 40% acid, you have a constant amount of acid and that is 20 ml So the acid fraction as you add water is 20/(50+x)

hero (hero):

I didn't even bother to calculate the acid fraction because she said the answer was in the back

hero (hero):

If you get a diagonal this time, check to see if you are set to radian (not degree) mode

OpenStudy (anonymous):

There's nothing shown on the graph.. o.o

hero (hero):

You're probably in degree mode. switch back to radians

OpenStudy (anonymous):

but It's on radians :(

hero (hero):

What kind of calculator do you have ?